OpenStudy (anonymous):

Evaluate f-^2g for f = 3 and g = 27.

OpenStudy (anonymous):

is it f^-2g or f^-2 x g?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f^{-2}\times g }\) \(\normalsize\color{blue}{ 3^{-2}\times 27 }\) \(\normalsize\color{blue}{ 3^{-2}\times 3^3 }\) \(\normalsize\color{blue}{ 3^{-2+3} }\) \(\normalsize\color{blue}{ 3^{?} }\)

OpenStudy (anonymous):

if it is the other way then  which is probably not it